OpenStudy (anonymous):

Suppose you have data that shows that 12% of athletes test positive for steroids. You also know that 11% of athletes test positive for steroids and actually use steroids. What is the probability that an athlete uses steroids, given that he tests positive? 0.37 0.43 0.51 0.67 0.92

OpenStudy (anonymous):

@HazelLuv99

OpenStudy (hazelluv99):

srry but i rlly can't help u i'm bad at math

OpenStudy (hazelluv99):

@Holly00d1248 PLEASE HELPPPPPPP

OpenStudy (anonymous):

lol u dont have to try so hard to help mee

OpenStudy (perl):

would you like help with this

OpenStudy (anonymous):

yeah

OpenStudy (holly00d1248):

No @perl he doesnt want help he just wants you to look at it

OpenStudy (perl):

lets label the probabilities

OpenStudy (perl):

P( test positive ) = .12 P ( use steroids& test positve ) = .11 P ( use steroids | test positive) = P ( use steroids & test positive ) / P( test positive)

OpenStudy (anonymous):

yes

OpenStudy (perl):

we are using the formula $$ \Large P( A| B ) = \frac{ P( A \& B ) }{P(B)} $$

OpenStudy (perl):

can you solve it now?

OpenStudy (anonymous):

ya i dont get it

OpenStudy (anonymous):

im about to guess on this i need to get it dun

OpenStudy (perl):

$$\Large \rm {P( use ~steroids | test~ positive)\\~\\ = \frac{ P( use ~steroids ~\&~ test~ positive ) }{P(test~ positive)} = \frac{.11}{.12} } $$

OpenStudy (anonymous):

.9166?

OpenStudy (perl):

you can round that

OpenStudy (anonymous):

thx for ur help got answer off goooooooooggggle

OpenStudy (perl):

can i see the google link

OpenStudy (anonymous):

can u help me with another question before i leave and yes openstudy.com/updates/51f93db8e4b06e6a116ea63c

OpenStudy (perl):

ok

OpenStudy (anonymous):

openstudy.com/updates/51f93db8e4b06e6a116ea63c

OpenStudy (perl):

google can give you the wrong answer, i wouldn't use google

OpenStudy (anonymous):

ohh

OpenStudy (perl):

http://prntscr.com/756jy6

OpenStudy (perl):

what did your google link answer you?